1st attempt See Periodic Table See Hint A sample of gas trapped by a moveable piston has a volume of 57.0L and a pressure of 2.70 atm at 273 K. If the volume of the gas decreases to 18.0L as the temperature is increased to 328 K, what is the new pressure of the gas? atm
Answer #1

PV = nRT

Since the gas is trapped, number of moles will be constant.

So, (P1*V1)/T1 = (P2*V2)/T2

(2.7*57.0)/273 = (P2*18.0)/328

P2 = 10.27 atm